OpenStudy (anonymous):

The area of a triangle must be at most 32.5sq in. The base is 13in and the height is x inches. Find the possible value of x. Use an inequality to solve the problem.

OpenStudy (anonymous):

formula for the area of a triangle is one half base times height. in symbols \(\frac{1}{2}bh\)

OpenStudy (anonymous):

in your case you have \(b=13, h = x\) and you want the area to be "at most" 32.5 meaning \[\frac{1}{2}\times 13x\leq 32.5\]

OpenStudy (anonymous):

to solve for x, write \[\frac{13}{2}x\leq 32.5\] \[x\geq 32.5\times \frac{2}{13}\] and then some arithmetic

OpenStudy (anonymous):

ok im lost